LIFTEZ LES SYLOWS! UNE SUITE À “SOUS-GROUPES PERIODIQUES D’UN GROUPE STABLE”

BRUNO POIZAT AND FRANK O WAGNER

Abstract. If G is an omega-stable group with a normal definable subgroup H, then the Sylow-2-subgroups of G/H are the images of the Sylow-2-subgroups of G.